# Quantum State Preparation via Free Binary Decision Diagram

### v2

@@ I Introduction
On the other hand, in many cases, we can employ suitable data structures to represent the coefficients of a quantum state in a compressed way, e.g., by using polynomial non-zero coefficients Zhang et al. (2022), discrete approximations of an efficiently integrable probability density function Grover and Rudolph (2002), a matrix product state (MPS) Holmes and Matsuura (2020); García-Ripoll (2021), a prior knowledge of the coefficients Sanders et al. (2019), and polynomial or Fourier approximations McArdle et al. (2022). In their settings, those results resolve the trade-off between the circuit depth, and the number of ancillary qubits for QSP. (See Table 1.)
@@
To address the challenge, we employ a data structure for the representation of Boolean functions, called free binary decision diagrams (FBDDs) Breitbar P et al. (1995). An $\mathrm{FBDD}_{f}$ to represent a Boolean function $f:x\in\{0,1\}^{n}\mapsto\{0,1\}$ is a rooted directed acyclic graph with one or two nodes of outdegree zero, and the other nodes of outdegree two, as illustrated in Fig. 1. We construct a quantum algorithm for QSP when the classical description of an $n$-qubit quantum state is given by an FBDD with weighted edges.
In this paper, we prove two claims. One is that any $\mathrm{FBDD}_{f}$ with $N$ nodes, and complex weighted edges represents a classical description of an $n$-qubit quantum state written by |$\ket{\psi_{f}}=\sum_{x\in\{0,1\}^{n}}c_{x}f(x)\ket{x}$, from which it is shown that the $\mathrm{FBDD}_{f}$ encodes a subset of the basis states designated by the Boolean function, i.e., the support {|$\{\ket{x}|\ x\in\{0,1\}^{n},f(x)=1\}$, and the weighted edges encode the coefficients $c_{x}$’s respectively. The other is that there exists a classical algorithm to construct a quantum circuit to prepare a quantum state represented by the weighted FBDD using $O(N)$ single-, and two-qubit gates, and $N$ ancillary qubits.
@@
We give a non-trivial example of exponentially compressible data structures by using a symmetric Boolean function Breitbar P et al. (1995). This result shows the existence of $n$-qubit quantum states with $d=O(2^{n})$ non-zero coefficients that can be prepared by a $\mathrm{poly}(n)$-sized quantum circuit with $\mathrm{poly}(n)$ ancillary qubits using our FBDD-based algorithm, but requires exponential amount of ancillary qubits using the algorithm for QSP in Ref. Zhang et al. (2022).
@@ II Preliminaries
Figure 3: FBDD, OBDD, and WFBDD for $f=\bar{x}_{1}\bar{x}_{2}x_{3}+\bar{x}_{1}x_{2}x_{4}+x_{1}\bar{x}_{3}x_{4}+x_{1}x_{2}x_{3}$. (a) In the FBDD, each path from the root to the terminal nodes has at most one occurrence of each variable. (b) In the OBDD, the order of occurrences of the variables on each path from the root to the terminal nodes is consistent. (c) In the WFBDD, weights are assigned to directed edges. When the head node of any directed edge is the 00-terminal node, the weight assigned to the directed edge is zero.
@@
This definition means that the WFBDD can be regarded as the BDD to which the contraction rules are applied first, and weights are assigned after the contraction. Note that a weighted $\mathrm{OBDD}$ ($\mathrm{WOBDD}$) is also defined by assigning non-zero weights to all directed edges of a given $\mathrm{OBDD}$ except ones of which the head nodes are the 00-terminal node.
@@ III.1 Classical description of quantum states by weighted FBDD
The path $P(z)$ specifies a sequence of nodes from the root to one of the terminal nodes. At each node $u$, the amplitude value is successively updated according to the ratio of two weights. The term $1/2^{n-|V_{z}|+1}$ of Eq. (2) comes from the deleted nodes having equal weighted edges. Let us show the above definition is well-defined, i.e., $|\alpha(z)|^{2}$ is a probability distribution.
@@
The following result (Theorem 7 of Ref. Breitbar P et al. (1995)) is known. For every $n\geq 4$, the $\mathrm{FBDD}_{h}$ is computable with no more than $33(2n^{2}+2n)+$ 3 nodes; however, every OBDD computing $h$ needs at least $2^{n/3}$ nodes. Thus, if the number of $z$ such as $h(z)=1$ is $O(poly(n))$, then the size of an OBDD computing $h$ is $O(poly(n))$, which contradicts Theorem 7 of Ref. Breitbar P et al. (1995). Therefore, the $\mathrm{FBDD}_{h}$ with weighted edges is a case of an $n$-qubit quantum state with $d=O(2^{n})$ non-zero coefficients.
@@ III.2 Quantum algorithm for QSP from weighted FBDD
First of all, we show that Algorithm 1 gives a stack $S$ of all internal nodes of $\textrm{WFBDD}_{f}$ with a linear total ordering such that for every directed edge $(u,v)$ out from an internal node $u$ to another internal node $v$, $u$ comes before $v$ in the ordering, i.e., a topological sort. In Algorithm 1, a set $Q=\{r\}$ consisting of the root $r$ is prepared (Line 4), $r$ is removed from $Q$ (Line 10), and pushed to $S$ (Line 11). Since $r$ is indegree 00, $S$ follows the ordering. Next, for each $b$-edge $e_{b}(r)$ out from $r$ to its head node $h_{b}(r)$, we remove $e_{b}(r)$ from $E$, and check the indegree of $h_{b}(r)$, and whether $h_{b}(r)$ is the terminal nodes $t_{0},t_{1}$ or not (Lines 14-16). If the indegree is zero, and $h_{b}(u)\neq t_{0},t_{1}$, then $h_{b}(r)$ is added to $Q$ (Line 17). The added node $h_{b}(r)$ is removed from $Q$, and pushed to $S$ in some round of the while loop. The head node $h_{b}(r)$ follows the ordering because the indegree zero means that all the directed edges whose head node is $h_{b}(r)$ have been removed from $E$, and the tail nodes of the directed edges have been pushed to $S$. Repeating this process, we order all the nodes. If there is an unordered internal node $u$, then the indegree of $u$ must be non-zero, which means another unordered internal node. Repeating this discussion leads that the root $r$ is not unordered, which contradicts that $r$ is first pushed to $S$. Let $\prec$ denote the total ordering.
@@
Assume that $\alpha_{u}(z)$, and $g_{u}(z)$ satisfy their definitions for an internal node $u$, and any internal node preceding to the node, and consider the internal node $v$ following the node $u$ in the ordering. For an arbitrary path $P(z)=(V(z),E(z))$ from the root $r$, through $v$, and to the terminal node $t_{1}$, let $e_{b}(u)=(u,v)\in E(z)$, and $e_{b^{\prime}}(v)=(v,w)\in E(z)$ denote directed edges into, and out from the node $v$, respectively. Since $g_{u}(z)$ satisfies its definition, and thus the qubit $v$ is set to |$\ket{1}$, applying the control-$H$ gate, and the control-$U(v)$ gate controlled by the qubit $v$ are applied to qubit $x_{i(v)}$ set to $\ket{+}$ (Lines 5, 12, and 13), we derive that
@@ IV Quantum state preparation for uniform superpositions
First, we show that Algorithm 2 gives a linear total ordering with which all the nodes follow. Through Algorithm 2, whether a node $u$ is ordered or not shall be determined by whether a set $X_{u}$ is empty or not. Thus, for each internal node $u$, $X_{u}$ is initialized to be an empty set, and for the terminal nodes $t_{0}$, and $t_{1}$, $X_{t_{0}}=X_{t_{1}}:=\{x_{1},\ldots,x_{n}\}$ (Lines 4, and 5). The terminal node $t_{1}$ is enqueued to a queue $Q$ (Line 7). For each tail node $u$ of all the directed edges into $t_{1}$, check whether both of $X_{h_{0}(u)}$, and $X_{h_{1}(u)}$ are non-empty or not (Lines 11, and 12), where $h_{b}(u)$ is the head node of the directed edge. If both of $X_{h_{0}(u)}$, and $X_{h_{1}(u)}$ are non-empty, $u$ is enqueued to $Q$, and $X_{u}$ is updated to be $X_{u}=(X_{h_{0}(u)}\cap X_{h_{1}(u)})\setminus\{x_{i(u)}\}$, where $i(u)$ is the index of the variable assigned to $u$ (Lines 12-14, and 18). Repeating this process, we order all the nodes, since if there exists a non-ordered node $u$, then at least either of $X_{h_{0}(u)}$ or $X_{h_{1}(u)}$ is empty which derives the contradiction that either of $X_{t_{0}}$ or $X_{t_{1}}$ is empty finally. Let $\prec$ denote the total ordering.
@@ VI Application to Black-Box QSP
In this section, we give an $n$-qubit quantum state that cannot be prepared efficiently by the black-box QSP Bausch (2022) but can be generated efficiently by our proposed QSP. For a given $\delta\in(0,1)$, we consider an $n$-qubit quantum state
@@ VII Application to the block encoding
The above two results can give non-trivial examples of block encodings of non-sparse matrices. For example, the block encodings of matrices derived from the $\mathrm{WFBDD}_{h}$ of which $h$ is the Boolean function of Eq. (12) are the cases, since the $\mathrm{WFBDD}_{h}$ represents a $n$-qubit quantum state with $d=O(2^{n})$ non-zero coefficients, as shown in Section III.
@@ VIII Conclusion
The efficiency of our QSP algorithm depends on the compression rate of WFBDDs. When $N=O(\mathrm{poly}(n))$, our algorithm gives an $O(\mathrm{poly}(n))$-sized QSP circuit with $\mathrm{poly}(n)$ ancillary qubits, which reduces the ancillary qubits needed in existing QSP algorithms for a general class of quantum states Zhang et al. (2022); Sun et al. (2023); Rosenthal (2021). Furthermore, we gave an example of $O(\mathrm{poly}(n))$-sized $\mathrm{WFBDD}_{f}$ representing an $n$-qubit quantum state with exponential non-zero coefficients. From these results, we contributed to propose a new subclass of efficiently preparable quantum states based on WFBDDs.
@@ Acknowledgements.
H.Y. was supported by JST PRESTO Grant Number JPMJPR201A, JPMJPR23FC, JSPS KAKENHI Grant Number JP23K19970, and MEXT Quantum Leap Flagship Program (MEXT QLEAP) JPMXS0118069605, JPMXS0120351339. M.M. was supported by MEXT QLEAP Q-LEAP JPMXS0118069605, and KAKENHI Grant No. 21H03394, 23K21643.

### v5

@@ I Introduction
[…]
@@
[…]
blackTo address the challenge, we employ a data structure for representing Boolean functions, called binary decision diagrams (BDDs) Breitbar P et al. (1995). A $\mathrm{BDD}_{f}$ representing a Boolean function $f:x\in\{0,1\}^{n}\mapsto\{0,1\}$ is a rooted directed acyclic graph with one or two nodes of outdegree zero and other nodes of outdegree two, where each node in the BDD represents a Boolean variable. In this paper, our method specifically utilizes a free binary decision diagram (FBDD), rather than a binary tree (BT) or an OBDD used in standard BDD-based QSP methods. QSP using FBDDs has not been explored because FBDDs do not share a key property that both BTs and OBDDs have. The property is that the order of occurrences of the Boolean variables on each path from the root to the terminal nodes is consistent. It assures that, at each depth, all the Boolean variables corresponding to nodes at the same depth represent the same qubit. Leveraging this property, for example, the method in Ref. Prakash (2014) achieves quantum state preparation (QSP) by coherently accessing the qubits labeled by nodes at the same depth.
In this paper, we construct a quantum algorithm for QSP when the classical description of an $n$-qubit quantum state is given by an FBDD with weighted edges. We prove two claims. One is that any $\mathrm{FBDD}_{f}$ with $N$ nodes, and complex weighted edges represents a classical description of an $n$-qubit quantum state written by |$\ket{\psi_{f}}=\sum_{x\in\{0,1\}^{n}}c_{x}f(x)\ket{x}$, from which it is shown that the $\mathrm{FBDD}_{f}$ encodes a subset of the basis states designated by the Boolean function, i.e., the support {|$\{\ket{x}|\ x\in\{0,1\}^{n},f(x)=1\}$, and the weighted edges encode the coefficients $c_{x}$’s respectively. The other is that there exists a classical algorithm to construct a quantum circuit to prepare a quantum state represented by the weighted FBDD using $O(N)$ single-, and two-qubit gates, and $N$ ancillary qubits.
@@
When $N=\mathrm{poly}(n)$, our FBDD-based algorithm gives a $\mathrm{poly}(n)$-sized quantum circuit for QSP with $\mathrm{poly}(n)$ ancillary qubits, which reduces the ancillary qubits needed for existing QSP algorithms for preparing a general class of states Zhang et al. (2022); Sun et al. (2023); Rosenthal (2021). Our FBDD-based algorithm exponentially reduces the number of two-qubit gates used by a QSP algorithm Mozafari et al. (2020) based on ordered BDDs (OBDDs), a proper subset of FBDDs Breitbar P et al. (1995).
@@
blackFurthermore, our FBDD-based algorithm also reduces exponentially the number of two-qubit gates used by another OBDD-based algorithm Mozafari et al. (2022) preparing an $n$-qubit quantum state described by an OBDD with $k=O(2^{n})$ paths. The approach in Ref. Mozafari et al. (2022) operates by preparing the paths in an OBDD one-by-one. As a result, the circuit complexity is linear in the number $k$ of paths in the OBDD, rather than in the number $N$ of nodes. It is important to emphasize that, even if $N$ is polynomial in the depth of the BDD, $k$ is not necessarily polynomial. In contrast, our method achieves circuit complexity that is linear in the number $N$ of nodes, which presents a significant advantage over the approach.
We give a non-trivial example of exponentially compressible data structures by using a symmetric Boolean function Breitbar P et al. (1995). This result shows the existence of $n$-qubit quantum states with $d=O(2^{n})$ nonzero coefficients that can be prepared by a $\mathrm{poly}(n)$-sized quantum circuit with $\mathrm{poly}(n)$ ancillary qubits using our FBDD-based algorithm, but requires exponential amount of ancillary qubits using the algorithm for QSP in Ref. Zhang et al. (2022).
[…]
blackReference Vinkhuijzen et al. (2024) provides an analytical investigation of three widely used classical descriptions of quantum states: MPS, decision diagrams (DDs), and restricted Boltzmann machines (RBMs). The authors map the relative succinctness of these data structures. Our results on weighted FBDDs are closely related to one of their results on the DDs called semiring-labeled decision diagrams (${\rm SLDD}_{\times}$), which is interpreted as weighted OBDDs in our paper’s terminology. This fact implies that the weighted FBDD can provide a more succinct classical description of quantum states than ${\rm SLDD}_{\times}$.
@@ II Preliminaries
Figure 3: FBDD, OBDD, and WFBDD for $f=\bar{x}_{1}\bar{x}_{2}x_{3}+\bar{x}_{1}x_{2}x_{4}+x_{1}\bar{x}_{3}x_{4}+x_{1}x_{2}x_{3}$. (a) In the FBDD, each path from the root to the terminal nodes has at most one occurrence of each variable. (b) In the OBDD, the order of occurrences of the variables on each path from the root to the terminal nodes is consistent. (c) In the WFBDD, directed edges are assigned weights. If the head node of a directed edge is the 00-terminal node, the weight of that edge is set to zero.
@@
This definition means that the WFBDD can be regarded as the BDD to which the contraction rules are applied first, and weights are assigned after the contraction. Note that a weighted $\mathrm{OBDD}$ ($\mathrm{WOBDD}$) is also defined by assigning nonzero weights to all directed edges of a given $\mathrm{OBDD}$ except ones of which the head nodes are the 00-terminal node.
@@ III.1 Classical description of quantum states by weighted FBDD
The path $P(z)$ specifies a sequence of nodes from the root to one of the terminal nodes. At each node $u$, the amplitude value is successively updated according to the ratio of two weights. The first multiplicative factor in Eq. (2) comes from the deleted nodes having equal weighted edges. Let us show the above definition is well-defined, i.e., $|\alpha(z)|^{2}$ is a probability distribution.
@@
The following result (Theorem 7 of Ref. Breitbar P et al. (1995)) is known. For every $n\geq 4$, the $\mathrm{FBDD}_{h}$ is computable with no more than $33(2n^{2}+2n)+$ 3 nodes; however, every OBDD computing $h$ needs at least $2^{n/3}$ nodes. Thus, if the number of $z$ such as $h(z)=1$ is $O(poly(n))$, then the size of an OBDD computing $h$ is $O(poly(n))$, which contradicts Theorem 7 of Ref. Breitbar P et al. (1995). Therefore, the $\mathrm{FBDD}_{h}$ with weighted edges is a case of an $n$-qubit quantum state with $d=O(2^{n})$ nonzero coefficients.
@@ III.2 Quantum algorithm for QSP from weighted FBDD
First of all, we show that Algorithm 1 gives a stack $S$ of all internal nodes of $\textrm{WFBDD}_{f}$ with a linear total ordering such that for every directed edge $(u,v)$ out from an internal node $u$ to another internal node $v$, $u$ comes before $v$ in the ordering, i.e., a topological sort. In Algorithm 1, a set $Q=\{r\}$ consisting of the root $r$ is prepared (Line 4), $r$ is removed from $Q$ (Line 10), and pushed to $S$ (Line 11). Since $r$ is indegree 00, $S$ follows the ordering. Next, for each $b$-edge $e_{b}(r)$ out from $r$ to its head node $h_{b}(r)$, we remove $e_{b}(r)$ from $E$, and check the indegree of $h_{b}(r)$, and whether $h_{b}(r)$ is the terminal nodes $t_{0},t_{1}$ or not (Lines 14-16). If the indegree is zero, and $h_{b}(u)\neq t_{0},t_{1}$, then $h_{b}(r)$ is added to $Q$ (Line 17). The added node $h_{b}(r)$ is removed from $Q$, and pushed to $S$ in some round of the while loop. The head node $h_{b}(r)$ follows the ordering because the indegree zero means that all the directed edges whose head node is $h_{b}(r)$ have been removed from $E$, and the tail nodes of the directed edges have been pushed to $S$. Repeating this process, we order all the nodes. If there is an unordered internal node $u$, then the indegree of $u$ must be nonzero, which means another unordered internal node. Repeating this discussion leads that the root $r$ is not unordered, which contradicts that $r$ is first pushed to $S$. Let $\prec$ denote the total ordering.
@@
Assume that $\alpha_{u}(z)$, and $g_{u}(z)$ satisfy their definitions for an internal node $u$, and any internal node preceding to the node, and consider the internal node $v$ following the node $u$ in the ordering. For an arbitrary path $P(z)=(V(z),E(z))$ from the root $r$, through $v$, and to the terminal node $t_{1}$, let $e_{b}(u)=(u,v)\in E(z)$, and $e_{b^{\prime}}(v)=(v,w)\in E(z)$ denote directed edges into, and out from the node $v$, respectively. Since $g_{u}(z)$ satisfies its definition and thus the qubit $v$ is set to |$\ket{1}$, Applying both of the controlled-$H$ gate and the controlled-$U(v)$ gate, controlled by the qubit $v$, to qubit $x_{i(v)}$ set to $\ket{+}$ (Lines 5, 12, and 13), we obtain that
@@ IV Quantum state preparation for uniform superpositions
First, we show that Algorithm 2 gives a linear total ordering with which all the nodes follow. In Algorithm 2, whether the set $X_{u}$ is empty serves as the criterion for determining the order of the node $u$. Thus, for each internal node $u$, $X_{u}$ is initialized to be an empty set, and for the terminal nodes $t_{0}$, and $t_{1}$, $X_{t_{0}}=X_{t_{1}}:=\{x_{1},\ldots,x_{n}\}$ (Lines 4, and 5). The terminal node $t_{1}$ is enqueued to a queue $Q$ (Line 7). For each tail node $u$ of all the directed edges into $t_{1}$, check whether both of $X_{h_{0}(u)}$, and $X_{h_{1}(u)}$ are non-empty or not (Lines 11, and 12), where $h_{b}(u)$ is the head node of the directed edge. If both of $X_{h_{0}(u)}$, and $X_{h_{1}(u)}$ are non-empty, $u$ is enqueued to $Q$, and $X_{u}$ is updated to be $X_{u}=(X_{h_{0}(u)}\cap X_{h_{1}(u)})\setminus\{x_{i(u)}\}$, where $i(u)$ is the index of the variable assigned to $u$ (Lines 12-14, and 18). Repeating this process, we order all the nodes, since if there exists a non-ordered node $u$, then at least either of $X_{h_{0}(u)}$ or $X_{h_{1}(u)}$ is empty which derives the contradiction that either of $X_{t_{0}}$ or $X_{t_{1}}$ is empty finally. Let $\prec$ denote the total ordering.
@@ VI Application to Black-Box QSP
In this section, we give an $n$-qubit quantum state that cannot be efficiently prepared by the black-box QSP Bausch (2022) but can be generated efficiently by our proposed QSP. For a given $\delta\in(0,1)$, we consider an $n$-qubit quantum state
@@ VII Application to the block encoding
The above two results can give non-trivial examples of block encodings of non-sparse matrices. For example, the block encodings of matrices derived from the $\mathrm{WFBDD}_{h}$ of which $h$ is the Boolean function of Eq. (12) are the cases, since the $\mathrm{WFBDD}_{h}$ represents a $n$-qubit quantum state with $d=O(2^{n})$ nonzero coefficients, as shown in Section III.
@@ VIII Conclusion
The efficiency of our QSP algorithm depends on the compression rate of WFBDDs. When $N=O(\mathrm{poly}(n))$, our algorithm gives an $O(\mathrm{poly}(n))$-sized QSP circuit with $\mathrm{poly}(n)$ ancillary qubits, which reduces the ancillary qubits needed in existing QSP algorithms for a general class of quantum states Zhang et al. (2022); Sun et al. (2023); Rosenthal (2021). Furthermore, we gave an example of $O(\mathrm{poly}(n))$-sized $\mathrm{WFBDD}_{f}$ representing an $n$-qubit quantum state with exponential nonzero coefficients. From these results, we contributed to propose a new subclass of efficiently preparable quantum states based on WFBDDs.
@@ Acknowledgements.
H.Y. was supported by JST PRESTO Grant Number JPMJPR201A, JPMJPR23FC, JSPS KAKENHI Grant Number JP23K19970, and MEXT Quantum Leap Flagship Program (MEXT QLEAP) JPMXS0118069605, JPMXS0120351339. M.M. was supported by MEXT QLEAP Q-LEAP JPMXS0118069605, and KAKENHI Grant No. 21H03394.